OpenStudy (anonymous):

Prove that log_2(3) is irrational. @ganeshie8 @myininaya @UnkleRhaukus @RadEn @robtobey @chmvijay @Luigi0210 @eliassaab @genius12 @Vincent-Lyon.Fr @Preetha @Kainui @ehuman @wolfe8 @lucaz @INeedHelpPlease? @khadeeja @thadyoung @A_clan @tester97 @Andras @Confusionist @linh412986 @BlackLabel @leozap1 @JMark @Microrobot @zacharyf @Rubio101 @NaomiBell1997 @MeganEdward @theballer225 @link,zoro @Goldenmoon @CrayolaCrayon_

OpenStudy (yttrium):

Assume log(base 2)3 is rational This means that log(base 2)3 = a/b where a and b are integers 2^(a/b) = 3 2^a = 3^b Note that 2^a MUST be even if 'a' is an integer Note that 3^b MUST be odd if 'b' is an integer Hence a and b cannot both be integers Hence a/b is irrational Hence, using proof by contradiction, log(base 2)3 is irrational

OpenStudy (anonymous):

To add to @Yttrium 's proof, we have that equality occurs when a = b = 0. And so it is better if we had initially defined that log_2(3)=a/b, where a and b are coprime positive integers. Since 2^a = 3^b is never true for positive integers a and b as the two sides are always of opposite parity, we reach a contradiction. Hence log_2(3) is irrational. @FutureMathProfessor